OpenStudy (anonymous):

I just began Logarithms and I'm a little confused... Can someone please help me? I'm posting the question in the comments.

OpenStudy (anonymous):

\[\log_{3} 243\]

OpenStudy (anonymous):

How is this read? Log of 3 and what?

OpenStudy (wolf1728):

Wow you are just beginning logarithms and already they give you a question about base 3 logs? I think that is read the base 3 log of 243

OpenStudy (anonymous):

okay, I wasn't sure

OpenStudy (wolf1728):

Now let's see how to calculate the base 3 log of 243 1 way to do it is with Excel (or even better the free online spreadsheet at openoffice.com

OpenStudy (wolf1728):

and the answer is 5

OpenStudy (wolf1728):

The base 3 log of 243 is 5 because 3^5 is 243.

OpenStudy (anonymous):

is there a way to write that out? Like to show my work.. Also, there's something about change of bases... I'm not sure how to do that either.

OpenStudy (wolf1728):

Not to change the subject but I think the easiest way to approach logs is with base 10 logs.

OpenStudy (anonymous):

what's that?

OpenStudy (wolf1728):

Okay well you just saw a base 3 log right?

OpenStudy (anonymous):

yeah

OpenStudy (wolf1728):

So, let's look at the powers of 10 10^1 = 10 10^2 = 100 10^3 = 1,000 10^4 = 10,000

OpenStudy (wolf1728):

Now if we wanted to know the base 10 logarithm of 100 it would be 2, because 2 is the exponent of 10^2

OpenStudy (anonymous):

I think the thing you're talking about is for this problem: \[\log_{b} 1000=3\]

OpenStudy (wolf1728):

That would be the base 10 log of 1,000 - yes

OpenStudy (anonymous):

okay, so it would be 3?

OpenStudy (anonymous):

I'm confused...

OpenStudy (wolf1728):

yes the base 10 log of 1,000 is 3

OpenStudy (anonymous):

what would b equal? The choices are 0, 10, 100, 1000

OpenStudy (wolf1728):

yes, I'm not surprised you are confused. Is this the first time you are learning about logarithms? I think you could find the answer to that question - want to try?

OpenStudy (anonymous):

yes, it is the first... I think it might be 100 butthat's just a guess

OpenStudy (anonymous):

oh wait....... Is it 10 because 10^3 = 1000?

OpenStudy (wolf1728):

The answer is right on this page. A few posts ago I said that the base 10 log of 1000 is 3 - and you got it right!!!

OpenStudy (anonymous):

Yay! by the way the explanation is written like this:  (these are the practice problems) Is that how it's done?